### v4

@@ A.2 Placebo Exercises
We use a placebo exercise to assess whether our choice of premium and non-premium investors is meaningful. To this end, we explore the effect of choosing an arbitrary set of PIPE investors and designating them as premium PIPE investors on the estimates. If our results were only noise, then a random group of premium PIPE investors should also give similar estimates. To implement this “placebo” exercise, we randomly picked 34 out of all 1,826 PIPE investors and labeled them premium investors. Then, we estimate all the models using this new set of premium investors. We repeat this exercise 500 times and present only the coefficients for premium and non-premium investors and the 95% confidence intervals in Figure A5, in the first and second columns, respectively. The five rows are redemption rate, announcement day return, mean-reversion, liquidation risk, and probability of liquidation.161616For the mean-reversion exercise, we present the first 500 cases with the first-stage F-stat .
@@
Variable |
Unit |
Definition |
@@
IPO proceeds |
MM USD |
Gross proceeds raised in the IPO, including any full or partial exercise of the Greenshoe. |
No. of Warrant |
Number of warrants included in the unit issued at the SPAC’s IPO. |
No. of Right |
Number of rights included in the unit issued at the SPAC’s IPO. |
Overallotment |
Amount of the overallotment option exercised by the underwriter(s) as a fraction of total IPO proceeds. |
Listing-day return |
SPAC IPO investors’ first-day return. |
Days searching |
Day |
The number of days between the SPAC’s IPO date and the target announcement date. |
Days remaining |
Day |
The number of days between the SPAC’s target announcement date and the liquidation deadline. |
Announcement-day return |
One-day return on the SPAC’s shares on the target’s announcement date. |
Redemption |
Redeemed SPAC common shares as a percentage of total shares issued at IPO. |
Days deSPAC |
Day |
The number of days between the SPAC’s target announcement date and the merger closing date. |
Cold |
pps |
The difference between the SPAC’s offer price and the SPAC’s share price on the closing day of the merger. |
Liquidation |
binary |
One if the SPAC liquidated during our sample period, zero otherwise. |
Target sector |
Sector which the SPAC’s registration statement indicates it will focus on in its search for a target for a business combination. |
Merger sector |
Sector in which the merger company operates. |
Jurisdiction |
State (if incorporated in the U.S.) or country in which the SPAC is incorporated at the time of the IPO. |
No. active SPACs |
Number of SPACs actively searching for targets at the time of the focal SPAC’s IPO. |
No. past successful deals |
Number of SPACs that completed the merger process before this focal SPAC’s IPO, and initiated by the same sponsor. |
@@
{PIPE (pre)} |
Binary |
One if PIPE (pre) investors participate in the deal. |
{PIPE (at)} |
Binary |
One if PIPE (at) investors participate in the deal. |
{PIPE (post)} |
Binary |
One if PIPE (post) investors participate in the deal. |
{Premium PIPE (at)} |
Binary |
One if premium PIPE (at) investors participate in the deal. |
{Non-Premium PIPE (at)} |
Binary |
One if non-premium PIPE (at) investors participate in the deal. |
No. of PIPE (at) |
Number of PIPE (at) investors. |
No. Premium PIPE (at) |
Number of premium PIPE (at) investors. |
No. of Non-Premium PIPE (at) |
Number of non-premium PIPE (at) investors. |
@@
%PIPE (pre) |
PIPE (pre) investors’ investment as a percentage of total IPO proceeds. |
%Premium PIPE (at) |
Premium PIPE (at) investors’ investment as a percentage of total IPO proceeds. |
%Non-Premium PIPE (at) |
Non-premium PIPE (at) investors’ investment as a percentage of total IPO proceeds. |
%PIPE (post) |
PIPE (post) investors’ investment as a percentage of total IPO proceeds. |
@@
Non-Premium |
Binary |
One if either the PIPE (at) investor’s number of participated PIPE deals or the total commitment amount is not in the top 5th percentile. |
Investment Amount |
MM USD |
Dollar amount of the PIPE investor’s investment in the SPAC deal. |
MLOT |
MM USD |
The difference between the value of securities held by the PIPE investor on the first day of listing of the post-merger company and the PIPE investor’s initial investment. |
lag-MLOT |
MM USD |
The PIPE investor’s MLOT from the previous SPAC deal with the same sponsor. |
lag-MLOT |
MM USD |
The PIPE investor’s MLOT from the one before the previous SPAC deal with the same sponsor. |
MLOT(past) |
MM USD |
Total MLOT a PIPE investor made in past deals with the sponsor. |
No. of past SPACs |
Number of SPAC deals participated by the PIPE investor before the focal SPAC deal. |
Past Investment Amount |
MM USD |
The PIPE investor’s total amount of investments in SPAC deals before the focal SPAC deal. |
No. of SPACs with other sponsors |
Number of other sponsors’ SPAC deals participated in by the PIPE investor before the focal SPAC deal. |
Past MLOT with other sponsors |
MM USD |
The PIPE investor’s total MLOT made in other sponsors’ SPACs before the focal SPAC deal. |
No. of SPACs with other sponsors, same industry |
Number of other sponsors’ SPAC deals that found targets in the same industry as the focal SPAC and participated by the PIPE investor before the focal SPAC deal. |
Past MLOT with other sponsors, same industry |
MM USD |
The PIPE investor’s total MLOT made in other sponsors’ SPACs that found targets in the same industry as the focal SPAC. |
Outside Option |
MM USD |
Highest MLOT among all premium PIPE investors not included in the current deal. |
Participation |
Binary |
One if the sponsor allocates securities to the non-premium PIPE investor, and zero otherwise. |
Allocation |
Ratio of the non-premium investor’s shares to the total shares allocated to the sponsor’s relationship investors. |

### v5

@@ A.2 Placebo Exercises
We use a placebo exercise to assess whether our choice of premium and non-premium investors is meaningful. To this end, we explore the effect of choosing an arbitrary set of PIPE investors and designating them as premium PIPE investors on the estimates. If our results were only noise, then a random group of premium PIPE investors should also give similar estimates. To implement this “placebo” exercise, we randomly picked 34 out of all 1,826 PIPE investors and labeled them premium investors. Then, we estimate all the models using this new set of premium investors. We repeat this exercise 500 times and present only the coefficients for premium and non-premium investors and the 95% confidence intervals in Figure A3, in the first and second columns, respectively. The five rows are redemption rate, announcement day return, mean-reversion, liquidation risk, and probability of liquidation.171717For the mean-reversion exercise, we present the first 500 cases with the first-stage F-stat .
@@
Appendix B Variable Definition
| Variable | Unit | Definition |
@@ Appendix B Variable Definition
| IPO proceeds | MM USD | Gross proceeds raised in the IPO, including any full or partial exercise of the Greenshoe. |
| No. of Warrant | Number of warrants included in the unit issued at the SPAC’s IPO. | |
| No. of Right | Number of rights included in the unit issued at the SPAC’s IPO. | |
| Overallotment | % | Amount of the overallotment option exercised by the underwriter(s) as a fraction of total IPO proceeds. |
| Listing-day return | % | SPAC IPO investors’ first-day return. |
| Days searching | Day | The number of days between the SPAC’s IPO date and the target announcement date. |
| Days remaining | Day | The number of days between the SPAC’s target announcement date and the liquidation deadline. |
| Announcement-day return | % | One-day return on the SPAC’s shares on the target’s announcement date. |
| Redemption | % | Redeemed SPAC common shares as a percentage of total shares issued at IPO. |
| Days deSPAC | Day | The number of days between the SPAC’s target announcement date and the merger closing date. |
| Cold | pps | The difference between the SPAC’s offer price and the SPAC’s share price on the closing day of the merger. |
| Liquidation | binary | One if the SPAC liquidated during our sample period, zero otherwise. |
| Target sector | Sector which the SPAC’s registration statement indicates it will focus on in its search for a target for a business combination. | |
| Merger sector | Sector in which the merger company operates. | |
| Jurisdiction | State (if incorporated in the U.S.) or country in which the SPAC is incorporated at the time of the IPO. | |
| No. active SPACs | Number of SPACs actively searching for targets at the time of the focal SPAC’s IPO. | |
| No. past successful deals | Number of SPACs that completed the merger process before this focal SPAC’s IPO, and initiated by the same sponsor. | |
@@
| {PIPE (pre)} | Binary | One if PIPE (pre) investors participate in the deal. |
| {PIPE (at)} | Binary | One if PIPE (at) investors participate in the deal. |
| {PIPE (post)} | Binary | One if PIPE (post) investors participate in the deal. |
| {Premium PIPE (at)} | Binary | One if premium PIPE (at) investors participate in the deal. |
| {Non-Premium PIPE (at)} | Binary | One if non-premium PIPE (at) investors participate in the deal. |
| No. of PIPE (at) | Number of PIPE (at) investors. | |
| No. Premium PIPE (at) | Number of premium PIPE (at) investors. | |
| No. of Non-Premium PIPE (at) | Number of non-premium PIPE (at) investors. | |
@@
| %PIPE (pre) | % | PIPE (pre) investors’ investment as a percentage of total IPO proceeds. |
| %Premium PIPE (at) | % | Premium PIPE (at) investors’ investment as a percentage of total IPO proceeds. |
| %Non-Premium PIPE (at) | % | Non-premium PIPE (at) investors’ investment as a percentage of total IPO proceeds. |
| %PIPE (post) | % | PIPE (post) investors’ investment as a percentage of total IPO proceeds. |
@@
| Non-Premium | Binary | One if either the PIPE (at) investor’s number of participated PIPE deals or the total commitment amount is not in the top 5th percentile. |
| Investment Amount | MM USD | Dollar amount of the PIPE investor’s investment in the SPAC deal. |
| MLOT | MM USD | The difference between the value of securities held by the PIPE investor on the first day of listing of the post-merger company and the PIPE investor’s initial investment. |
| lag-MLOT | MM USD | The PIPE investor’s MLOT from the previous SPAC deal with the same sponsor. |
| lag2-MLOT | MM USD | The PIPE investor’s MLOT from the one before the previous SPAC deal with the same sponsor. |
| MLOT(past) | MM USD | Total MLOT a PIPE investor made in past deals with the sponsor. |
| No. of past SPACs | Number of SPAC deals participated by the PIPE investor before the focal SPAC deal. | |
| Past Investment Amount | MM USD | The PIPE investor’s total amount of investments in SPAC deals before the focal SPAC deal. |
| No. of SPACs with other sponsors | Number of other sponsors’ SPAC deals participated in by the PIPE investor before the focal SPAC deal. | |
| Past MLOT with other sponsors | MM USD | The PIPE investor’s total MLOT made in other sponsors’ SPACs before the focal SPAC deal. |
| No. of SPACs with other sponsors, same industry | Number of other sponsors’ SPAC deals that found targets in the same industry as the focal SPAC and participated by the PIPE investor before the focal SPAC deal. | |
| Past MLOT with other sponsors, same industry | MM USD | The PIPE investor’s total MLOT made in other sponsors’ SPACs that found targets in the same industry as the focal SPAC. |
| Outside Option | MM USD | Highest MLOT among all premium PIPE investors not included in the current deal. |
| Participation | Binary | One if the sponsor allocates securities to the non-premium PIPE investor, and zero otherwise. |
| Allocation | Ratio of the non-premium investor’s shares to the total shares allocated to the sponsor’s relationship investors. | |